Joseph is 7 years older than his brother Tony.Twice Joseph's age in 4 years minus 2 will equal three times Tony's age in 4 years How old is joseph now ?

Answer by stanbon(75887)

Joseph is 7 years older than his brother Tony.Twice Joseph's age in 4 years minus 2 will equal three times Tony's age in 4 years How old is joseph now ?
------------
Equations:
j = t + 7
2(j+4) -2 = 3(t+4)
-----
Substitute for "j" and solve for "t"::
2(t+11) -2 = 3t + 12
-----
2t + 20 = 3t + 12
t = 8 yrs (Tony's age now)
-----
j = t+7 = 15 yrs (Joe's age now)
-------------
cheers,
Stan H.
-----------